need to find the x,y intercepts of this equation
4x-6y = -24
thank you
dave

Answer by drglass(89)

There are a few ways to do this, perhaps the easiest is to find the value of y when x = 0 (the y-intercept) and the value of x when y = 0 (the x-intercept).


First, the y intercept:
4(0) -6y = -24 or
-6y = -24 or
y = 4


The y-intercept is (0, 4)


Next, the x-intercept:
4x -6(0) = -24 or
4x = -24 or
x = -6


The x-intercept is (-6, 0)